OpenStudy (anonymous):

\[\int_0^{2\pi}\int_0^{pi/4}\int_0^4(\rho\cos\phi)\rho^2\sin\phi \ \ d\rho \ \ d\phi \ \ d\theta \]

OpenStudy (anonymous):

:-( I should have picked medicine over engineering

OpenStudy (anonymous):

first compute this by treating fi as a constant: \[\int_0^4(\rho\cos\phi)\rho^2\sin\phi \ \ d\rho\]

OpenStudy (anonymous):

so sin(phi) is taken out?

OpenStudy (anonymous):

\[\int_0^4(\rho\cos\phi)\rho^2\sin\phi \ \ d\rho=sin\phi cos\phi\int_0^4\rho^3 \ d\rho\] if I read it correctly